OpenStudy (anonymous):

Simplify by factoring. Assume that all expressions under radicals represent nonnegative numbers. ^3sqrt108x^8

OpenStudy (anonymous):

\[\sqrt[3]{108x^8}\] \[\sqrt[3]{3^32^2x^3x^3x^2}\] \[3x^2\sqrt[3]{4x^2}\]